OpenStudy (anonymous):

Solve 3x + 2y = –12 for y

OpenStudy (anonymous):

@Valkarie70

OpenStudy (anonymous):

ok, how many questions do you still have? if there is multiple you can put them down 5 at a time. also i hope you are learning not just me giving them to you, watch what i do in my work because i wont always be here to answer them. 3x+2y=-12 -3x -3x 2y=-12-3x 2y=-3x-12 to look nicer /2 /2 Answer: y=-1.5x-6

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

??

OpenStudy (anonymous):

choices are y=3/2x-6 y= -3/2x-6 y=3/2x-12 y=-3/2x-12

OpenStudy (anonymous):

@Valkarie70

OpenStudy (anonymous):

owe then its B y=-3/2x-6 BTW thats the same thing as what I wrote if you looked at my work, and next time plz put in the awnswers